# An Uncertainty Quantification Framework for Deep Learning-Based Automatic Modulation Classification

### v1

@@ V Conclusion
Deep Learning (DL) has been shown to provide cutting-edge performance in automatic modification classification (AMC). However, DL-based AMC models often exhibit overconfidence in their predictions, with no associated measure of uncertainty. This issue is especially evident in low signal to noise ratio (SNR) conditions and when handling out-of-distribution (OOD) samples. In this work, we proposed a deep ensemble framework for AMC, which is capable of retaining the state-of-the-art performance of DL-based AMC classifiers while simultaneously providing robust uncertainty quantification (UQ) metrics. We demonstrated our ensemble’s ability to achieve robust performance across multiple UQ metrics such as the negative log-likelihood, Brier score, prediction interval widths, prediction coverage, and high-confidence predictions. In comparison to standalone CNNs as well as Bayesian Neural Networks (BNNs), we showed that our framework achieved better UQ estimates overall, particularly in low SNR and OOD environments. Moreover, our proposed framework is scalable to any DL architecture, allowing state-of-the-art performance to be extended to incorporate higher classification performance as well as uncertainty quantification on any DL-based AMC framework. Future work will explore the resilience of our approach in more versatile environments such as the distributed AMC scenario, which requires UQ in federated learning with varying channel conditions and adversarial interference at each receiver, and in environments with insufficient channel state information (CSI) knowledge.

### v2

@@ V Conclusion
Deep learning (DL) has been shown to provide cutting-edge performance in automatic modulation classification (AMC). However, DL-based AMC models often exhibit overconfidence in their predictions, with no associated measure of uncertainty. This issue is especially evident in low signal to noise ratio (SNR) conditions and when handling out-of-distribution (OOD) samples. In this work, we proposed a deep ensemble framework for AMC, which is capable of retaining the state-of-the-art performance of DL-based AMC classifiers while simultaneously providing robust uncertainty quantification (UQ) metrics. We demonstrated our ensemble’s ability to achieve robust performance across multiple UQ metrics such as the negative log-likelihood, Brier score, prediction interval widths, prediction coverage, and high-confidence predictions. In comparison to standalone CNNs, weighted ensembles, and Bayesian Neural Networks (BNNs), we showed that our framework achieved better UQ estimates overall, particularly in low SNR and OOD environments. Moreover, our proposed framework is scalable to any DL architecture, allowing state-of-the-art performance to be extended to incorporate higher classification performance as well as uncertainty quantification on any DL-based AMC framework. Future work will explore the resilience of our approach in more versatile environments such as the distributed AMC scenario, which requires UQ in federated learning with varying channel conditions and adversarial interference at each receiver, and in environments with insufficient channel state information (CSI) knowledge.
[Uncaptioned image] Huian Yang is currently pursuing his B.S. in Computer Engineering at the University of California San Diego. His research focuses on deep learning, with additional interests in embedded systems for robotics.
[Uncaptioned image] Rajeev Sahay received the B.S. degree in electrical engineering from The University of Utah, Salt Lake City, UT, USA, in 2018, and the M.S. and Ph.D. degrees in electrical and computer engineering from Purdue University, West Lafayette, IN, USA, in 2021 and 2022, respectively. Currently, he is a faculty member in the Department of Electrical and Computer Engineering at UC San Diego. He was the recipient of the Purdue Engineering Dean’s Teaching Fellowship and was named an Exemplary Reviewer by the IEEE Wireless Communications Letters. His research interests lie in the intersection of networking and machine learning, especially in their applications to wireless communications and engineering education.